# Compress then Serve: Serving Thousands of LoRA Adapters with Little Overhead

### v1

@@ 5.1 Training
We trained LoRA adapters on 500 natural instruction tasks [27] using Mistral-7B-Instruct-v0.2 [10] as the base model. All LoRA adapters were configured with a rank of 16, i.e., $\forall i,r_{i}=16$.
We selected 10 diverse tasks (Table 2 in Appendix D) manually for consistent evaluation across experiments and randomly sampled an additional 490 tasks, resulting in a total of 500 tasks. These tasks were exclusively in English (both input and output), ensuring higher quality and thorough review [27]. Each task dataset was divided into training, validation, and test sets.
Hyperparameters, such as early stopping, were tuned using the validation sets. Evaluation on the test sets demonstrated that LoRA consistently outperformed the base model in terms of both Rouge scores and loss metrics, as shown in Table 1. Details are provided in Appendix D.
@@ 5.2 Evaluation
We evaluated multiple metrics for the natural instruction tasks, including cross-entropy loss, Rouge-1, Rouge-L [15], exact match, and agreement between uncompressed and compressed LoRA. Here, agreement measures the exact match in task-generations between the uncompressed LoRA model and the compressed LoRA model, rather than comparing to ground truth data. While detailed results and discussions for all metrics are provided in Appendix F, our primary focus in the main text is on Rouge-L. We find that all metrics correlate, but Rouge-L correlates most strongly with downstream utility. This finding aligns with prior work [27], which demonstrates that Rouge-L correlates well with classification accuracy.
@@ 6.1 Task Performance
Figure 2(a) plots the mean performance improvement relative to uncompressed LoRAs ($y$-axis) against the GPU workload parameter saved ratio, i.e., the reduction in per-LoRA parameters ($x$-axis). Similarly, Figure 2(b) relates mean performance improvement to the total parameter saved ratio. Both the independent compression via SVD and joint diagonalization methods significantly compress the LoRAs while preserving—and occasionally enhancing—performance. Notably, the JD methods are unique in their ability to approach the compression efficacy of a single LoRA, although this aggressive reduction in size may decrease performance in larger LoRA collections. As shown in Figure 4, a JD variant achieves the highest throughput. These results provide practitioners with a spectrum of methods and configurations, enabling them to optimize the balance between throughput, compression, and performance according to their specific requirements.
For efficiency, we limited the JD methods to ten iterations instead of pursuing full convergence. While the alternating algorithm quickly reaches an approximation of the minimizer, squeezing out the last few digits of precision takes many more iterations with limited to no performance gain. Appendix F.12 also evaluates an alternative eigenvalue iteration algorithm that more rapidly converges once $U,V$ are close to a minimizer of the optimization problem, with minimal performance differences.
(a) Performance with GPU parameter saved ratio
(b) Performance with total parameter saved rate
Figure 2: We qualitatively illustrate the performance of JD-LoRA. The GPU workload parameter saved ratio indicates the number of parameters saved compared to LoRA during inference when switching between different LoRAs. For JD-Diagonal methods, it is computed as $r_{pergpu}:=1-\frac{r_{JD}}{r_{LoRA}*d*2}$. The total parameter saved ratio depicts the number of parameters saved for a system with a large number $K$ of LoRAs for different tasks. It is computed as: $r_{total}:=1-\frac{2*r_{JD}*d+K*r_{JD}}{K*r_{LoRA}*d*2}$, where $r_{JD}$ and $r_{LoRA}$ are the rank for JD-LoRA and LoRA respectively, $d$ is the weight dimension, and $K$ is the number of LoRAs served.
@@ 6.2 Performance and Reconstruction Error
Figure 4 relates reconstruction error and performance. The $y$-axis measures mean performance improvement of Rouge-L relative to uncompressed LoRA, and the $x$-axis quantifies the mean reconstruction error between the compressed reconstruction of the product $BA$ and the original uncompressed product $BA$. Although the relationship between performance and reconstruction error is nonlinear, it demonstrates a generally decreasing, somewhat exponential trend. Notably, the minimal reconstruction error does not correlate with optimal performance, indicating that a degree of lossy reconstruction may be advantageous for enhancing generalization.
@@
Figure 4: Throughput when serving varying number of LoRAs with vLLM.
@@ 6.3 Throughput of Serving Compressed LoRAs
Figure 4 studies the efficacy of compression in a real-world serving scenario. We consider a varying number of rank-16 LoRAs, using a dataset of Shakespeare sonnets as inputs [^2] https://www.kaggle.com/datasets/shivamshinde123/william-shakespeares-sonnet/data arriving asynchronously. We measure throughput, i.e., the number of requests served per second when generating ten tokens per request.
In these experiments, the base model is Mistral 7B Instruct as in the previous experiments; we simulate random LoRAs and assign inputs to LoRAs at random. Experiments were conducted on A100 80GB GPU capped at 30% memory consumption. This is done to reflect cost concerns in practical situations where a service provider might want to serve many LoRAs from cheaper hardware with lower memory than higher-end GPUs. This setting also takes into account the scenario where the LLM is large compared to the size of GPU and yet a provider may want to serve many LoRAs efficiently using the same device.
Although LoRAs are meant as parameter-efficient adapters, when serving over 200 LoRAs, the throughput starts to degrade rapidly, illustrating the challenge of serving many LoRAs. When serving over 500 LoRAs, our compression methods provide over 2×2$\times$2 × improvement in throughput, and at over 1000 LoRAs, JD-Full-64 serves at 75% of the throughput of the base LLM (equivalent to serving a single LoRA after merging it with the LLM).
Next, we analyze trade-offs between the compression methods. Each compression method except JD-Full-64 and JD-Diag-128 starts to degrade faster at a certain point as the number of LoRAs increases; this point is the threshold at which compressed LoRAs no longer fit into the GPU memory and the throughput degrades due to scheduling. The JD LoRAs are the most efficient in terms of memory consumption, and we can simultaneously load all 1024 LoRAs. Due to their increased rank (e.g., each LoRA is of rank 128 for JD-Diag-128), however, the throughput starts lower, but degrades at a lower rate. We conclude that for the current vLLM design, when serving a moderate number of LoRAs, SVD compression provides favorable throughput as long as it can fit all LoRAs into memory. When serving a large number of LoRAs that cannot fit into memory even with SVD compression, JD is the better option, as it can fit all LoRAs onto the GPU at the cost of a slight throughput reduction due to its larger rank.
Our analysis above suggests that scheduling LoRAs is an important bottleneck. Compressed LoRAs, especially with JD, significantly reduce the amount of parameters that need to be loaded onto a GPU to process a request requiring an offloaded LoRA. Designing a custom scheduler that can leverage this property is a promising avenue to further improve the throughput when serving compressed LoRAs.
@@
We also note that vLLM extensively uses custom CUDA kernels. To accommodate our compression techniques, we minimally adjusted the vLLM code to generate additional kernels needed by the compressed LoRAs. There likely is room for improvement to optimize the newly added kernels.
Fast LoRA [28] is a recent work that aims to alleviate the batched matrix multiplication (BMM) bottleneck when serving many LoRAs. They propose an adapter parameterization that replaces addition with elementwise multiplication, avoiding BMM and improving LoRA throughput at lower ranks. Our JD LoRA formulation also circumvents or heavily reduces the impact of BMM (see Appendix E), and both individual and joint compression methods can be applied to Fast LoRAs.
@@ 7 Discussion
This study introduces approaches to LoRA compression, addressing significant challenges facing foundation models and large language models. Our contributions include theoretical formulations, empirical validation, and practical implementations that enhance understanding and application of LLMs in scalable environments.
@@
A primary limitation involves integration of new LoRAs. Our findings indicate that while compressing a new LoRA within the original space improves upon the baseline, it can degrades performance compared to uncompressed LoRAs (see Appendix F.11). Thus, the optimal strategy varies with the frequency of new additions. Infrequent additions justify recompressing all LoRAs given the minimal impact relative to inference costs. Conversely, in scenarios where a proportion of new LoRAs are often introduced but seldom served, one should compress only frequently-served LoRAs and otherwise use the original LoRAs.
The promising results of our study suggest several future research directions. Extending our compression techniques to a wider array of models and tasks could further substantiate and expand the generalizability of our findings. This includes out-of-distribution evaluations, which can highlight generalization improvements associated with compression. Moreover, more sophisticated optimization algorithms could improve the balance between compression and performance.
@@
8 Acknowledgements
The MIT Geometric Data Processing Group acknowledges the generous support of Army Research Office grants W911NF2010168 and W911NF2110293, from the CSAIL Systems that Learn program, from the MIT–IBM Watson AI Laboratory, from the Toyota–CSAIL Joint Research Center, and from an Amazon Research Award.
Appendix A Limitations
Compression techniques, while reducing time or memory constraints, invariably lose information, which can increase with the number of models compressed. Consequently, while our approach seeks to minimize this loss, it is axiomatic that compression will diminish information to some degree.
Our methodology compresses a collection of known models. Introducing new models may necessitate either recompression, projecting into a suboptimal basis, or opting to use some models in an uncompressed state.
Empirically, our methods achieve reasonable balances between the number of models compressed, the amount of information retained, and throughput. Nonetheless, these benefits can be accompanied by increased latency as more models are compressed.
Appendix B Joint Diagonalization Algorithms
B.1 Alternating Methods
Our goal is to derive algorithms that optimize (2). Common to both methods, we expand the objective functional:
@@ B.1 Alternating Methods
Case 1: Non-diagonal $\Sigma_{i}$, orthogonal $U,V$. Setting the derivative of (5) with respect to $\Sigma_{i}$ to zero, we find
@@
by the identity $\|A\|_{\mathrm{Fro}}^{2}=\mathrm{tr}(A^{\top}A)$. Hence, we optimize (7) by alternating between $U$ and $V$:
@@
Case 2: Diagonal $\Sigma_{i}$. If constrain $\Sigma_{i}$ to be diagonal, we interpret our objective function (2) as a “triple least squares” problem. We compute gradients:
@@
$=(U^{\top}B_{i}\circ V^{\top}A_{i}^{\top})\mathbbm{1}$
$\implies\mathrm{diag}(\Sigma_{i})$ $=(U^{\top}U\circ V^{\top}V)^{-1}(U^{\top}B_{i}\circ V^{\top}A_{i}^{\top})\mathbbm{1}$
@@
B.2 Additional Eigenvalue Iteration Algorithm
For the first case in §B.1, we introduce an alternative algorithm that eschews the use of SVD. This alternative is optimized for GPU execution, enabling tractable runs to convergence.
To derive this algorithm, we employ Lagrange multipliers to formulate the derived objective from (7):
@@ B.2 Additional Eigenvalue Iteration Algorithm
Here, one can show that the Lagrange multiplier matrices $X$ and $Y$ are diagonal and nonnegative, since the problem reduces to an eigenvalue problem when either $U$ or $V$ is fixed; this is essentially the argument behind the alternating algorithm in Appendix B. Hence, taking inspiration from classical eigenvalue iteration, we use the following updates to improve our estimates of $U$ and $V$:
@@
Appendix C Proof of Theorem 1
@@
By the definition of singular value decomposition, the right hand side of (18) is maximized with $U$, $V$ being the top $r$ singular vectors of $\sum_{i=1}^{n}B_{i}A_{i}$, yielding $\left\|U^{\top}\sum_{i=1}^{n}B_{i}A_{i}V\right\|_{\mathrm{Fro}}^{2}=\sum_{i=1}^{r}\bar{\sigma}^{2}_{i}$. Recalling that $\Sigma_{i}=U^{\top}B_{i}A_{i}V$ yields the lower bound.
@@
Appendix D Training LoRAs
We trained LoRA adapters on 500 natural instruction tasks [27] using Mistral-7B-Instruct-v0.2 [10] as the base model. All LoRA adapters were configured with a rank of 16, i.e., $\forall i,r_{i}=16$. We selected 10 diverse tasks manually for consistent evaluation across experiments and randomly sampled an additional 490 tasks, resulting in a total of 500 tasks. These tasks were exclusively in English (both input and output), ensuring higher quality and thorough review [27]. Each task dataset was divided into training, validation, and test sets (80-10-10). Hyperparameters, such as early stopping, were tuned using the validation sets; that is, we train for five epochs and take the best-performing epoch-checkpoint per validation loss. Evaluation on the test sets demonstrated that LoRA consistently outperformed the base model in terms of both Rouge scores and loss metrics (see Table 1).
@@ Appendix D Training LoRAs
We use Huggingface [29] in our implementation. For the base model, we use quantization with configuration:
@@
Appendix E Avoiding Batched Matrix Multiplication (BMM)

### v2

@@ 5.1 Training
We trained LoRA adapters on 500 natural instruction tasks (Wang et al., 2022) using Mistral-7B-Instruct-v0.2 (Jiang et al., 2023a) as the base model. All LoRA adapters were configured with a rank of 16, i.e., $\forall i,r_{i}=16$.
We selected 10 diverse tasks (Table 2 in Appendix C) manually for consistent evaluation across experiments and randomly sampled an additional 490 tasks, resulting in a total of 500 tasks. These tasks were exclusively in English (both input and output), ensuring higher quality and thorough review (Wang et al., 2022). Each task dataset was divided into training, validation, and test sets.
Hyperparameters, such as early stopping, were tuned using the validation sets. Evaluation on the test sets demonstrated that LoRA consistently outperformed the base model in terms of both Rouge scores and loss metrics, as shown in Table 1. Details are provided in Appendix C.
@@ 5.2 Evaluation
We evaluated multiple metrics for the natural instruction tasks, including cross-entropy loss, Rouge-1, Rouge-L (Lin, 2004), exact match, and agreement between uncompressed and compressed LoRA. Here, agreement measures the exact match in task-generations between the uncompressed LoRA model and the compressed LoRA model, rather than comparing to ground truth data. While detailed results and discussions for all metrics are provided in Appendix G, our primary focus in the main text is on Rouge-L. We find that all metrics correlate, but Rouge-L correlates most strongly with downstream utility. This finding aligns with prior work (Wang et al., 2022), which demonstrates that Rouge-L correlates well with classification accuracy.
@@ 6.1 Task Performance
Figure 2 illustrates the Total Parameter Saved Ratio versus the Number of Unique LoRAs served. We only include methods that maintain over 99% of the original LoRA’s performance (as measured by RougeL). Notably, our JD methods uniquely approach the compression efficacy of a single LoRA, and with clustering, this aggressive reduction in size also maintains performance in larger LoRA collections. It is also interesting to note (see Appendix G) that JD variants often increase generalization and outperform the original LoRA.
For efficiency, we limited the JD methods to ten iterations instead of pursuing full convergence. While the alternating algorithm quickly reaches an approximation of the minimizer, squeezing out the last few digits of precision takes many more iterations with limited to no performance gain. Appendix G.10 also evaluates an alternative eigenvalue iteration algorithm that more rapidly converges once $U,V$ are close to a minimizer, with minimal performance differences.
Figure 2: Total Parameter Saved Ratio versus the Number of Unique LoRAs served. We only include methods that maintain over 99% of the original LoRA’s performance (as measured by RougeL). The Total Parameter Saved Ratio depicts the number of parameters saved for a system with a large number $n$ of different LoRAs. It is computed as: $r_{total}:=1-\frac{\text{num. parameters after compression}}{\text{num. parameters before compression}}$.
@@ 6.2 Performance and Reconstruction Error
Figure 4 relates reconstruction error and performance. The $y$-axis measures the mean performance improvement of Rouge-L relative to uncompressed LoRA, and the $x$-axis quantifies the mean reconstruction error between the compressed reconstruction of the product $BA$ and the original uncompressed product $BA$. Although the relationship between performance and reconstruction error is nonlinear, it demonstrates a generally decreasing, somewhat exponential trend. Notably, the minimal reconstruction error does not correlate with optimal performance, indicating that a degree of lossy reconstruction may be advantageous for enhancing generalization.
To select hyperparameters (compression rank and number of clusters) for the clustering experiments, we first assessed reconstruction error on a single LoRA module over a range of settings (see Appendix F). These preliminary experiments enabled efficient selection of cluster counts and rank values for compressing all LoRA modules.
@@
Figure 4: Throughput ratio when serving varying numbers of LoRAs with vLLM.
@@ 6.3 Throughput of Serving Compressed LoRAs
Results in the previous sections demonstrate how to select an appropriate joint compression setting guided by the reconstruction error, such that the performance of the original LoRAs is preserved. Naturally, the rank and/or the number of clusters for the compression needs to increase as we compress larger LoRA collections to match LoRA performance.
In Figure 4 we study how throughput with various compression settings compares to the vLLM multi-LoRA throughput with the matched GPU memory footprint. Specifically, for each number of unique LoRAs served and each compression setting, we compute the corresponding number of LoRAs to be placed on the GPU during serving and report the ratio of the two throughputs. For example, when serving 64 unique LoRAs and using rank 64 JD-Full compression, we report the ratio of throughputs of rank 64 JD-Full and vLLM multi-LoRA with 6 LoRAs allowed on the GPU at a time (see Appendix E for details). As the number of unique LoRAs increases, vLLM multi-LoRA throughput degrades as it needs to schedule the requests and load and offload the adapters. We note that vLLM multi-LoRA already employs many advanced system optimization techniques, such as efficient scheduling and non-blocking CPU-GPU communication when swapping LoRAs (Sheng et al., 2023; Kwon et al., 2023), but system optimization alone is not sufficient to fully mitigate throughput degradation when serving many LoRAs.
In Figure 4 we see that across all LoRA collection sizes our compression techniques improve the throughput of vLLM multi-LoRA. Additionally, we highlight regions for each compression setting where compression is sufficiently moderate to achieve 99%+ of LoRA performance, according to the results in Section 6.2. We also note that compression with a larger rank or too many clusters does not improve baseline throughput when serving a smaller number of LoRAs and should not be used in such cases. For example, rank 16 JD-Full improves baseline throughput with 4 and 8 LoRAs, but will underperform with more LoRAs, while 7 cluster rank 64 JD-Full does not improve throughput with 64 or fewer LoRAs, but when serving 1000+ LoRAs it improves the throughput significantly while maintaining the performance. To conclude, an appropriate joint compression setting improves vLLM multi-LoRA throughput and preserves performance for LoRA collections of any size between 4 and 1024, as we showed in Figure 1. Specific compression settings for each LoRA collection size are listed in Appendix E.
@@
Finally, we note that vLLM extensively uses custom CUDA kernels. To accommodate our compression techniques, we minimally adjusted the vLLM code to generate additional kernels needed by the compressed LoRAs. There likely is room for improvement to optimize the newly added kernels.
Additional details
In this experiment we considered a varying number of rank-16 LoRAs, using a dataset of Shakespeare sonnets as inputs [^2] https://www.kaggle.com/datasets/shivamshinde123/william-shakespeares-sonnet/data arriving asynchronously. We measured throughput, i.e., the number of requests served per second when generating ten tokens per request. The base model was Mistral 7B Instruct as in the other experiments; we simulated random LoRAs and assigned inputs to LoRAs at random. Experiments were conducted on H100 80GB GPU capped at 40% memory consumption. This was done to reflect cost concerns in practical situations where a service provider might want to serve many LoRAs from cheaper hardware with lower memory than higher-end GPUs. This setting also takes into account the scenario where the LLM is large compared to the size of GPU and yet a provider may want to serve many LoRAs efficiently using the same device.
@@ 7 Discussion
This study introduces approaches to LoRA compression, addressing significant challenges facing foundation models and large language models. Our contributions include theoretical formulations, empirical validation, and practical implementations that enhance the understanding and application of LLMs in scalable environments.
@@
The promising results of our study suggest several future research directions. First, further compression may be possible via quantization. Our joint-diagonalization compression and quantization are independent axes of approaching the problem and exploring a combined solution can be fruitful. Second, when scaling to hundreds of thousands of LoRAs, joint compression, while effective, will not be sufficient to fit all LoRAs onto the GPU, thus requiring a procedure to schedule the requests. Our clustering variant offers opportunities to develop an efficient scheduling mechanism that takes into account the cluster assignments of LoRAs corresponding to the incoming requests.
@@
Appendix A Joint Diagonalization Algorithms
A.1 Alternating Methods
Our goal is to derive algorithms that optimize equation 1. Common to both methods, we expand the objective functional:
@@ A.1 Alternating Methods
Case 1: Non-diagonal $\Sigma_{i}$, orthogonal $U,V$. Setting the derivative of equation 5 with respect to $\Sigma_{i}$ to zero, we find
@@
by the identity $\|A\|_{\mathrm{Fro}}^{2}=\mathrm{tr}(A^{\top}A)$. Hence, we optimize equation 7 by alternating between $U$ and $V$:
@@
Case 2: Diagonal $\Sigma_{i}$. If constrain $\Sigma_{i}$ to be diagonal, we interpret our objective function equation 1 as a “triple least squares” problem. We compute gradients:
@@
$=(U^{\top}B_{i}\circ V^{\top}A_{i}^{\top})\bm{1}$
$\implies\mathrm{diag}(\Sigma_{i})$ $=(U^{\top}U\circ V^{\top}V)^{-1}(U^{\top}B_{i}\circ V^{\top}A_{i}^{\top})\bm{1}$
@@
A.2 Additional Eigenvalue Iteration Algorithm
For the first case in §A.1, we introduce an alternative algorithm that eschews the use of SVD. This alternative is optimized for GPU execution, enabling tractable runs to convergence.
To derive this algorithm, we employ Lagrange multipliers to formulate the derived objective from equation 7:
@@ A.2 Additional Eigenvalue Iteration Algorithm
Here, one can show that the Lagrange multiplier matrices $X$ and $Y$ are diagonal and nonnegative, since the problem reduces to an eigenvalue problem when either $U$ or $V$ is fixed; this is essentially the argument behind the alternating algorithm in Appendix A. Hence, taking inspiration from classical eigenvalue iteration, we use the following updates to improve our estimates of $U$ and $V$:
@@
A.3 Clustering algorithm
Initialization: We run joint diagonalization with a single $U,V$ then perform k-means with $|C_{j}|$ clusters on the space of $\Sigma_{i}$’s. This gives us our first clusters and we can use random initialization $U_{j},V_{j}$ for each cluster but the $\Sigma_{i}$ can be maintained as initialization.
Step 1: Using the alternating JD algorithms from earlier in this section, we optimize the problem $\min_{U_{j},V_{j},\Sigma_{i}}\sum_{i\in C_{j}}||B_{i}A_{i}-U_{j}\Sigma_{i}V_{j}^{\top}||^{2}_{F}$ for each $j$ independently.
Step 2: New cluster assignment for $i$ : $\min_{j}\min_{\Sigma_{i}}||B_{i}A_{i}-U_{j}\Sigma_{i}V_{j}^{\top}||^{2}_{F}$. If any assignment changes we go to Step 1, else we have converged.
Appendix B Proof of Theorem 1
@@
By the definition of singular value decomposition, the right hand side of equation 18 is maximized with $U$, $V$ being the top $r$ singular vectors of $\sum_{i=1}^{n}B_{i}A_{i}$, yielding $\left\|U^{\top}\sum_{i=1}^{n}B_{i}A_{i}V\right\|_{\mathrm{Fro}}^{2}=\sum_{i=1}^{r}\bar{\sigma}^{2}_{i}$. Recalling that $\Sigma_{i}=U^{\top}B_{i}A_{i}V$ yields the lower bound.
@@
Appendix C Training LoRAs
We trained LoRA adapters on 500 natural instruction tasks (Wang et al., 2022) using Mistral-7B-Instruct-v0.2 (Jiang et al., 2023a) as the base model. All LoRA adapters were configured with a rank of 16, i.e., $\forall i,r_{i}=16$. We selected 10 diverse tasks manually for consistent evaluation across experiments and randomly sampled an additional 490 tasks, resulting in a total of 500 tasks. These tasks were exclusively in English (both input and output), ensuring higher quality and thorough review (Wang et al., 2022). Each task dataset was divided into training, validation, and test sets (80-10-10). Hyperparameters, such as early stopping, were tuned using the validation sets; that is, we train for five epochs and take the best-performing epoch-checkpoint per validation loss. Evaluation on the test sets demonstrated that LoRA consistently outperformed the base model in terms of both Rouge scores and loss metrics (see Table 1).
@@ Appendix C Training LoRAs
We use Huggingface (Wolf et al., 2020) in our implementation. For the base model, we use quantization with configuration:
@@
Appendix D Avoiding Batched Matrix Multiplication (BMM)
Fast LoRA (Wen & Chaudhuri, 2024) aims to alleviate the batched matrix multiplication (BMM) bottleneck when serving many LoRAs. They propose an adapter parameterization that replaces addition with elementwise multiplication, avoiding BMM and improving LoRA throughput at lower ranks. Our JD LoRA formulation also circumvents or heavily reduces the impact of BMM as discussed below, and both individual and joint compression methods can be applied to Fast LoRAs.